Find out about the wall construction (number and type of layers) of a sauna (treatment and/or relaxation room), thereby determining the heat transfer mode, the range of ambient conditions under which they operate, typical heat transfer coefficients on the inner and outer surfaces of the wall, and the heat generation rates inside. Determine the size and methodology of the heating and air-conditioning system that will be able to maintain the typical sauna conditions, for when 10 persons are inside the sauna at the same time.


1
Expert's answer
2022-01-05T22:51:01-0500

Sauna are rooms which are kept between "65°c \>\>to\>\>90°c" with a humidity rate of "5\%" to "10\%" .


They are meant to provide relaxation and varying types of health benefits.



Sauna ceilings, walls and floor are made from wood and are moisture resistant.

The interior of the sauna is usually covered with wood panelling for saunas and there is heat insulation between them and the wall. Mineral wool is used most oftenly used.

Most wall thickness are "1-\frac{3}{8}''" to "1-\frac{1}{2}''" and "1-\frac{5}{8}''"

The thicker the wall the greater the insulation.




In a hot air sauna, the user is heated through convection.

In a "steam bath" the water is boiled and the resulting steam heats the body through direct contact with the skin.

The emerging infrared sauna uses electromagnetic radiation from infrared lamps to warm body directly.



A sauna is heated between "60°C\>to\> 80°C" with a relative humidity of "5\%-30\%" .

Saunas have a vent to bring fresh air and release the humid air keeping the humidity of the sauna low.

They have several seat levels and the higher the seat the higher the temperature.


The Batholoey Acoustic 430 speakers are specifically for use in saunas and can withstand the extreme temperature and humidity

Sauna luminaires with LED lighting are perfectly safe, economic and reliable.





A sauna have a convective heat transfer coefficient of around "hi=10000W\/m^2.K"

Typical heat transfer outside "h_0=15W\/m^2.K"


A Large sauna that can accommodate 10 persons is powered by a 8kW Harvia KIP Electric sauna Heater.

A simple to operate dial control allow for easy access and comfortable control of heater with adjustable temperature range between "(158-194)\u00b0F"


Apart from Natural Ventilation, a mechanical Ventilation going into the domain HVAC (Heating, Ventilation, air conditioning) should be used in 10 people Sauna to ensure air circulates inside sauna through complicated ducts and heat exchangers.